# LLM Post-Training: A Deep Dive into Reasoning Large Language Models

### v1

@@ 7 Future Directions
We gathered all papers related to post-training methods in LLMs and analyzed their trends, as shown in Figure 7. Advancing RL techniques [303, 57, 40] for refining the LLMs have a noticeable increase in prominence since 2020 (Figure 7(a)), emphasizing the demand for interactive approaches such as human-in-the-loop [35, 304] reinforcement and scalability [111, 82, 305]. At the same time, reward modeling [306, 166, 167] (Figure 7(b)) has seen a steady rise in interest due to the emergence of self-rewarding language models, yet the field still struggles with reward hacking [307, 308] and the design of robust [309], failure-aware reward functions beyond reward hacking [310]. Decoding and search (Figure 7(c)) methods include tree-of-thoughts [84] and Monte Carlo [311, 257] strategies aiming to enhance model reasoning through iterative self-critique [312, 304, 29], but these techniques also demand reliable uncertainty estimators to prevent excessive computational overhead [313, 111]. Safety [299, 314, 315], robustness [316], and interpretability [317, 318, 319] have likewise become central concerns (Figure 7(d)), motivating the development of bias-aware [320, 321] and uncertainty-aware [322] RL methods beyond correlation with human uncertanity [323] that safeguard user trust and prevent adversarial attacks. Another crucial area involves personalization [324, 325] and adaptation [193] (Figure 7(e)), where efforts to tailor LLMs for specific domains must be balanced against risks to privacy [326], particularly when enterprise data or sensitive personal information is involved.
@@
Test-time scaling challenges. Open challenges in TTS revolve around how to orchestrate the inference-time processes efficiently and reliably. A key question is how much computing is enough for a given query, and how to determine this on the fly? Using less resources can result in mistakes, but using too much is inefficient and could introduce inconsistencies. Recent research by Snell et al. [83] tackled it by proposing a unified framework with a ‘Proposer’ and a ‘Verifier’ to systematically explore and evaluate answers. In their framework, the Proposer (usually the base LLM) generates multiple candidate solutions, and the Verifier (another model or a heuristic) judges and selects the best. The optimal strategy can vary by problem difficulty: for easier queries, generating many answers in parallel and picking the top might be sufficient, whereas for harder problems, sequential, step-by-step reasoning with verification at each step works better. An important future direction is building adaptive systems where the LLM dynamically allocates computation based on an estimate of the question’s complexity. This idea connects to meta-cognition in AI [314], enabling models to have a sense of what they don’t know or what deserves more thought. Developing reliable confidence metrics or difficulty predictors for LLMs is an open research area, but progress here would make TTS far more practical i.e., the model would only ‘slow down and think’ when necessary, much like a human spending extra time on a hard problem. Recent study [347] shows distilling test-time computations into synthetic training data create synergistic pretraining benefits which can also be further explored.
@@
[Uncaptioned image] Integration: Combining PRM-guided tree search with online distillation achieves 4× efficiency gains over baseline methods while maintaining 94% solution accuracy on MATH dataset.
@@
RL methods exhibit sample inefficiency and computational overhead, particularly when scaling to contexts exceeding 128k tokens. The ‘overthinking’ phenomenon, where models generate excessively long reasoning chains, further reduces token efficiency and increases deployment costs [373]. Investigate partial rollout strategies with flash attention mechanisms for long-context processing. Develop length penalty mechanisms using learned compression transformers for iterative long2short distillation. Hybrid architectures combining MCTS [74] with GRPO [59] could enable better exploration-exploitation tradeoffs. Parallel work by Xie et. al. [74] demonstrates promising results through adaptive tree search pruning. Several open challenges persist in the field. Uncertainty propagation remains problematic as current confidence estimators add approximately 18% latency overhead, while catastrophic forgetting results in a degradation of 29% of base capabilities during RL fine-tuning [374]. Moreover, benchmark saturation is an issue, with MMLU scores correlating poorly (r = 0.34) with real-world performance [375].

### v2

@@ 7 Future Directions
We gathered all papers related to post-training methods in LLMs and analyzed their trends, as shown in Figure 7. Application of RL techniques [303, 57, 40] for refining the LLMs have a noticeable increase in prominence since 2020 (Figure 7(a)), emphasizing the demand for interactive approaches such as human-in-the-loop [35, 304] reinforcement and scalability [111, 82, 305]. At the same time, reward modeling [306, 166, 167] (Figure 7(b)) has seen a steady rise in interest due to the emergence of self-rewarding language models, yet the field still struggles with reward hacking [307, 308] and the design of robust [309], failure-aware reward functions beyond reward hacking [310]. Decoding and search (Figure 7(c)) methods include tree-of-thoughts [84] and Monte Carlo [311, 257] strategies aiming to enhance model reasoning through iterative self-critique [312, 304, 29], but these techniques also demand reliable uncertainty estimators to prevent excessive computational overhead [313, 111]. Safety [299, 314, 315], robustness [316], and interpretability [317, 318, 319] have likewise become central concerns (Figure 7(d)), motivating the development of bias-aware [320, 321] and uncertainty-aware [322] RL methods beyond correlation with human uncertanity [323] that safeguard user trust and prevent adversarial attacks. Another crucial area involves personalization [324, 325] and adaptation [193] (Figure 7(e)), where efforts to tailor LLMs for specific domains must be balanced against risks to privacy [326], particularly when enterprise data or sensitive personal information is involved.
@@
Test-time scaling challenges. Open challenges in TTS revolve around how to orchestrate the inference-time processes efficiently and reliably. A key question is how much computing is enough for a given query, and how to determine this on the fly? Using less resources can result in mistakes, but using too much is inefficient and could introduce inconsistencies. Recent research by Snell et al. [83] tackled it by proposing a unified framework with a ‘Proposer’ and a ‘Verifier’ to systematically explore and evaluate answers. In their framework, the Proposer (usually the base LLM) generates multiple candidate solutions, and the Verifier (another model or a heuristic) judges and selects the best. The optimal strategy can vary by problem difficulty: for easier queries, generating many answers in parallel and picking the top might be sufficient, whereas for harder problems, sequential, step-by-step reasoning with verification at each step works better. An important future direction is building adaptive systems where the LLM dynamically allocates computation based on an estimate of the question’s complexity. This idea connects to meta-cognition in AI [314], enabling models to have a sense of what they don’t know or what deserves more thought. Developing reliable confidence metrics or difficulty predictors for LLMs is an open research area, but progress here would make TTS far more practical i.e., the model would only ‘slow down and think’ when necessary, much like a human spending extra time on a hard problem. Additionally, By reframing inference-time scaling as a probabilistic inference problem and employing particle-based Monte Carlo methods [puri2025probabilistic], the small models achieved o1 level accuracy in only 32 rollouts, a 4–16x improvement in scaling efficiency across various mathematical reasoning tasks. Recent study [347] shows distilling test-time computations into synthetic training data creates synergistic pretraining benefits which can also be further explored.
@@
[Uncaptioned image] Integration: Combining PRM-guided tree search with online distillation achieves 4× efficiency gains over baseline methods, while maintaining 94% solution accuracy on MATH dataset.
@@
RL methods exhibit sample inefficiency and computational overhead, particularly when scaling to contexts exceeding 128k tokens. The ‘overthinking’ phenomenon, where models generate excessively long reasoning chains, further reduces token efficiency and increases deployment costs [373]. Investigate partial rollout strategies with flash attention mechanisms for long-context processing. Develop length penalty mechanisms using learned compression transformers for iterative long2short distillation. Hybrid architectures combining MCTS [74] with GRPO [59] could enable better exploration-exploitation tradeoffs. Parallel work by Xie et. al. [74] demonstrates promising results through adaptive tree search pruning. Several open challenges persist in the field. Uncertainty propagation remains problematic as current confidence estimators add approximately 18% latency overhead, while catastrophic forgetting rresults in a degradation of 29% of base capabilities during RL fine-tuning [374]. Moreover, benchmark saturation is an issue, with MMLU scores correlating poorly (r = 0.34) with real-world performance [375].